OpenStudy (anonymous):

Simplify: 12x^5y^8/ 4x^8y^5

OpenStudy (anonymous):

Can you write it with parenthesis please? It's hard to know whether there's a multiplication or not

OpenStudy (phi):

you can simplify the numbers: 12 divided by 4 is what ? x^5 is short for x*x*x*x*x (x times itself 5 times) if you divide by x^8, x*x*x*x*x*x*x*x each x divided by x becomes 1. \[ \frac{ \cancel{x} \cdot \cancel{x} \cdot \cancel{x} \cdot \cancel{x} \cdot \cancel{x} } {\cancel{x} \cdot \cancel{x} \cdot \cancel{x} \cdot \cancel{x} \cdot \cancel{x} \cdot x \cdot x \cdot x }\]

OpenStudy (anonymous):

(12x^5y^8) / (4x^8y^5)

OpenStudy (anonymous):

3.

OpenStudy (phi):

yes, 12/4 is 3. so 3 is part of the answer.

OpenStudy (phi):

can you simplify the y's ? \[ \frac{y^8}{y^5} \] notice you have y*y*y*y*y in the bottom and they will "cancel" 5 of the y's up top

OpenStudy (anonymous):

shouldnt it be y^3 than ?

OpenStudy (phi):

yes you have y*y*y left over "up top" we can use short-hand and say y^3 so now you have \[ \frac{3 y^3 x^5}{x^8} \] if you simplify the x's you get x^3 in the bottom

OpenStudy (anonymous):

now what do I have to do

OpenStudy (anonymous):

simplify that ?

OpenStudy (phi):

simplify the x's

OpenStudy (anonymous):

1/x^3

OpenStudy (phi):

so the whole answer is \[ \frac{3y^3}{x^3} \] there is a "rule" that says that \[ \frac{1}{x^n} = x^{-n} \] so you could also write the answer as \[ 3 x^{-3}y^3 \]

OpenStudy (anonymous):

nice thank you somuch